OpenStudy (anonymous):

A country's population in 1995 was 228 million. In 2001 it was 230 million. Estimate the population in 2010 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (dumbcow):

233.03294 geez you really need to learn how to use a calculator to compute equations or how will you answer these on your own

OpenStudy (anonymous):

well if i knew how to i would be able to but seeing i don't know how to figure out the equation, it's a little hard to use a calculator to figure out the equation.

OpenStudy (dumbcow):

no i understand that, thats why i showed you which formula to use and gave you the steps but you have been able to compute the number from there

OpenStudy (anonymous):

like the one you gave me had e or something like that & idk what e is so i can't figure it out if i don't know

OpenStudy (dumbcow):

ok but that is something you should know to be taking this class all of your questions are going to be involving stuff like that go to your calculator and find the "e" and "ln" learn how to use them

OpenStudy (anonymous):

i don't have a graphing calculator & i do online schooling so i don't always learn everything i need cause the tutorial don't give you alot of detail

OpenStudy (dumbcow):

e is just a special number, about 2.71 ln is the inverse to e^x

OpenStudy (dumbcow):

any scientific calculator will have them download software that will compute things for you www.graphcalc.com

OpenStudy (anonymous):

do you think you could help me with my next problem

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

A country's population in 1991 was 136 million. In 2000 it was 141 million. Estimate the population in 2016 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (dumbcow):

do you want the steps, or just a number

OpenStudy (anonymous):

either

OpenStudy (dumbcow):

ok well steps would be the same just different numbers k = ln(141/136)/9 P25 = 136*e^((25/9)*ln(141/136)) 25/9 = 2.77 ln(141/136) = .0361 e^(2.77*.0361) = 1.10517 P25 = 136*1.10517 P25 = 150.3

OpenStudy (anonymous):

thank you!

OpenStudy (dumbcow):

try using calculator to get same answers i did